What is the price of a $1000 par value bond with a coupon rate of 11.3% (annual payments) that matures in 22 years, assuming that the bond’s YTM is 4.66% p.a.?
Answer #1

Solving problem with manual formula and calculator both:

#

I/Y = R = Rate or yield / frequency of coupon in a year =

                    4.660000

PMT = Coupon rate x FV / frequency =

$113.00

N = Number of years remaining x frequency =

22.00

FV = Face value =

1000

Formula for bond value:

PV = PMT x ((1-((1+R%)^-N)) / R%) + (FV/(1+R%)^N)

PV = (113* ((1-(1+0.0466)^-22)/0.0466) + 1000/(1+0.0466)^22) =

$1,901.77

=======================

Using financial calculator BA II Plus - Input details:

#

I/Y = R = Rate or yield / frequency of coupon in a year =

                    4.660000

PMT = Coupon rate x FV / frequency =

-$113.00

N = Number of years remaining x frequency =

22.00

FV = Future Value =

-$1,000.00

CPT > PV = Present value of bond = Price of Bond = Current value of bond =

$1,901.77